During 2016, Linda, Inc.’s net income was $1,400,000. Its common stockholders’ equity was $2,160,000 at January 1, 2016 and $3,440,000 at December 31, 2016. During 2016, Cooper had 10,000 outstanding shares of 6%, $300 par value cumulative preferred stock. During December, 2016, Linda’s board of directors declared the annual preferred stock dividend and a $240,000 common stock dividend.

What is Linda’s 2016 return on common stockholders’ equity?

Select one:

A. 50.0%

B. 41.4%

C. 43.6%

D. 35.0%
Answer #1

Return on common stockholders equity =

( net income) - (preference dividend ) / average common stock equity

Average common stock equity =( 2160000+3440000 ) / 2

= 2800000

Preference dividend =( 300*10000)*6% = 180000

Return on common stock holders equity =

((1400000 - 180000)/ 2800000 )*100

= 43.5714

= 43.6 %